OpenStudy (anonymous):

PLZ HELP WILL FAN AND MEDAL WHICH OF THE FOLLOWING HAS THE LOWEST UNIT PRICE? BRAND A: 5 IB FOR 3.50$ BRAND B: 8 IB FOR 5.20$ BRAND C: 3 IB FOR 2.25$ A. BRAND A B. BRAND B C. BRAND C D. ALL BRANDS HAVE THE SAME UNIT PRICE

OpenStudy (alexandervonhumboldt2):

lets find first the unit prices: BRAND A: 5 IB FOR 3.50$ unit price = 3.50/5=0.70 BRAND B: 8 IB FOR 5.20$ 5.20/8=0.65 BRAND C: 3 IB FOR 2.25$ 2.25/3=0.75

OpenStudy (anonymous):

WHAT DO YOU MEAN I THINK THE ANSWER I B

OpenStudy (alexandervonhumboldt2):

could you stop using caps? i'm not blind B is correct

OpenStudy (anonymous):

thanks and sorry I had my caps on because I like it on

OpenStudy (alexandervonhumboldt2):

ok